Derive Planck’s law of black body radiation.


1
Expert's answer
2021-03-12T07:14:45-0500

Answer

The probability that a single mode

has energy En = nhν is given by the usual Boltzmann factor

"P(n) =\frac{exp(-\frac{h\nu }{kT}) }{\sum exp(-\frac{h\nu }{kT}) }"


Let x="{exp(-\frac{h\nu }{kT}) }=\frac{h\nu }{kT}" For classicla limit

So average energy

"U=\sum E P(n)"


"U=kT"

The energy density of radiation range is

"u(\nu) d\nu =\frac{8\pi v^2 U d\nu}{c^3}\\ =\frac{8\pi hv^3 d\nu}{c^3(e^(\frac{h\nu}{kT}) -1)}"

This is planck radiation distribution law.